C_{2v}-TOP INTERNAL ROTATION IN THE HIGH-RESOLUTION UV SPECTRA OF WEAKLY BOUND COMPLEXES: II. APPLICATIONS.^{a}

\underline{MARTIN SCHÄFER} and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

The model of a C_{2v} top – C_{s} frame internal rotation discussed in Part I has been used in the analysis of rotationally resolved S_{1} ← S_{0} fluorescence excitation spectra of three different complexes; aniline–N_{2}, para-difluorobenzene–N_{2}, and benzonitrile–H_{2}O. The results of the internal rotation analysis as well as the structure determination will be presented and the limits of the applied model will be discussed.

aWork supported by the Swiss National Science Foundation and NSF.